OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2

OpenStudy (anonymous):

A. -sin^2Θ B. -cos^2Θ c. 2 d. 0

OpenStudy (anonymous):

please please help

OpenStudy (kropot72):

The first step is to expand the brackets. Expanding the first bracket we get \[\large (\sin\theta-\cos\theta)^{2}=\cos^{2}\theta-2\sin\theta \cos\theta+\cos^{2}\theta\] What do you get when you expand the second bracket?

OpenStudy (anonymous):

hm not sure

OpenStudy (anonymous):

please help me out real quick

OpenStudy (kropot72):

The expansion of the second bracket is \[\large \sin^{2}\theta+\sin\theta \cos\theta+\cos^{2}\theta\] Now you need to collect like terms in the sum of the two expansions and simplify.

OpenStudy (anonymous):

okay i understand so would the answer be A?

OpenStudy (kropot72):

Please don't guess the answer. The sum of the two expansions is \[\large 2\sin^{2}\theta+2\cos^{2}\theta\] Now you need to simplify.

OpenStudy (anonymous):

so how would i simplify this??

OpenStudy (kropot72):

Well 2 is a common factor. So we get \[\large 2(\sin^{2}\theta+\cos^{2}\theta)\] If you know your trig identities this can be easily simplified.

OpenStudy (kropot72):

What is the value of \[\large \sin^{2}\theta+\cos^{2}\theta=?\]

OpenStudy (anonymous):

im not sure im terrible at this

OpenStudy (anonymous):

would it equal 0 maybe?

OpenStudy (kropot72):

Not really. \[\large \sin^{2}\theta+\cos^{2}\theta=1\]

OpenStudy (anonymous):

so then you woiuld do 1 multiplyed by 2 and the answer would be 2?

OpenStudy (kropot72):

Correct!

OpenStudy (anonymous):

thank you so much

OpenStudy (kropot72):

You're welcome :)